1 A person deposited EUR 200,000 he was holding in cash into a bank account. Determine how the money supply in a banking system with 100%

reserves changes.

Select one from the following options.

i.The money supply will increase by EUR 200,000 times a money multiplier.

ii. The money supply will increase by EUR 200,000.

iii. The money supply will not change.

iv. The money supply will increase by EUR 100,000.


2 Consider the life cycle hypothesis. A consumer's consumption function in the form C = 0.025W + 0.5Y implies that: Select one from the following options.

i. all offered options are correct.

ii. consumption will increase by EUR 25, when wealth will increase by EUR 1,000.

iii. life expectancy is 40 years

iv. Half of a person's life expectancy is spent in retirement




 


1
Expert's answer
2021-05-31T09:09:54-0400

Solution:

1.). The correct answer is: ii. The money supply will increase by EUR 200,000.


Change in Money Supply = "Change\; in\; Reserves\times Money\; Multiplier"

Money Multiplier = "\frac{1}{Reserve \; Ratio} = \frac{1}{100\%} =1"


Change in Money Supply = 200,000 "\times" 1 = 200,000


The money supply will increase by EUR 200,000.


2.). The correct answer is: iii. life expectancy is 40 years